OpenStudy (anonymous):

Sides of three square rooms measure 13 feet each, and sides of two square rooms measure 15 feet each. Which expression shows the total area of these five rooms?

OpenStudy (anonymous):

So for the first 3 squares with sides of 13 : 13*13=165 165*3 = 495 15*15=225 225*2 = 450 495+450=945